OpenStudy (anonymous):

The admission fee of a theater is $2.50 for adults and $1.25 for children. On a certain day, 700 people went to the theater for a concert and $1375 was collected. How many children and how many adults attended the concert? a. 300 adults, 400 children b. 400 adults, 1100 children c. 400 adults, 300 children d. 600 adults, 100 children

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

2.50x + 1.25y = 1375 x + y = 700 Solve the system of equations. That's how you solve this problem.

OpenStudy (anonymous):

a + c = 700 Elimination 2.50a + 1.25c = 1375 -2.50( a + c = 700 )  Multiply everything in top equ. by –2.50 2.50a + 1.25c = 1375 -2.50a – 2.50c = -1750 2.50a + 1.25c = 1375 -1.25c = -375 -1.25c = -375 -1.25 -1.25 a + (300) = 700 c = 300 a = 400

OpenStudy (anonymous):

A.) i got the same answer